OpenStudy (kmullis6):

Choose the equation of the line passing through the point (-3, 1) and parallel to y = -x - 4. y = x + 2 y = -x - 2 y = -x + 4 y = x - 4

OpenStudy (haseeb96):

The slope of the given line is m=-1 Now the condition of parallel then slopes will be same.m=-1 (x,1,y1)=(-3,1) Now just these things into this formula y-y1=m(x,x1) And show me your work after putting the values...

OpenStudy (kmullis6):

I don't get how to do this.

OpenStudy (haseeb96):

y-1=-1(x+3) y-1=-x-3 y=-x-3+1 y=-x-2 Is your answer.

OpenStudy (haseeb96):

If u like my explanation then u can click on best response

OpenStudy (kmullis6):

Ok, thank you so much!

OpenStudy (haseeb96):

Welcome my sister.... U r cute